OpenStudy (anonymous):

SOLVE FOR X (1/3X)+6

OpenStudy (anonymous):

19/3

OpenStudy (anonymous):

multiply by x =(1/3)+6 Find LCD (3) 6*(3/3)+(1/3) =(18/3)+(1/3) (18+1)/3 Final answer: 19/3 =)